OpenStudy (anonymous):

Simplify: x^2-20x+14 ----------- x+3 what is the integer remainder?

OpenStudy (amistre64):

x -23 -------------- x+3 | x^2 -20x +14 -3x --------- -23x +14 +69 ----- ?? <- remainder

OpenStudy (anonymous):

for solving this substitue x=-3 in the equation x^2-20+14 and u'll find it...

OpenStudy (amistre64):

-3 | 1 -20 14 0 -3 +69 ----------- 1 -23 ?? <- remainder :)

OpenStudy (anonymous):

the question actually would be , wat is the remainder wen this quadratic equation is divided by x+3, which she wrote like this...

OpenStudy (anonymous):

P(x)=x^2-20x+14 P(-3)=9+60+14=83

OpenStudy (anonymous):

what uzma said. no reason to divide if you just want the remainder

OpenStudy (amistre64):

no reason to divide? lol

OpenStudy (anonymous):

woah so whose do i go by?

OpenStudy (anonymous):

yea, use remainder theorem

OpenStudy (amistre64):

well, only one person actually gave you the answer in the end ....

OpenStudy (amistre64):

personally, I was hoping you had enough sense about you to be able to add 14 and 69 ....

OpenStudy (anonymous):

yes amistre is right because question said "simplify" a terrible word but i guess in this case it means to actually divide

OpenStudy (anonymous):

so if it says "divide" you had best divide!

OpenStudy (amistre64):

yeah, I wanted to factor at first lol

OpenStudy (anonymous):

because you didn't finish reading the question, just like i didn't read the first part.

OpenStudy (anonymous):

so the integer remainder is 83?

OpenStudy (anonymous):

yes. amsitre gave you the answer. it is \[\frac{x^2+20x-14}{x+3}=x-23+\frac{83}{x+3}\]

OpenStudy (amistre64):

say it with confidence, it sounds better; the integer remainder is 83! :)

OpenStudy (anonymous):

which is another way of saying \[x^2+20x-14=(x+3)(x-23)+83\]

OpenStudy (anonymous):

thanks guys.

OpenStudy (amistre64):

:) youre welcome